### 27 Feb 2024 v1

# NocPlace: Nocturnal Visual Place Recognition Using Generative and Inherited Knowledge Transfer

Abstract
Visual Place Recognition (VPR) is crucial in computer vision, aiming to retrieve database images similar to a query image from an extensive collection of known images. However, like many vision-related tasks, learning-based VPR often experiences a decline in performance during nighttime due to the scarcity of nighttime images. Specifically, VPR needs to address the cross-domain problem of night-to-day rather than just the issue of a single nighttime domain. In response to these issues, we present NocPlace, which leverages a generated large-scale, multi-view, nighttime VPR dataset to embed resilience against dazzling lights and extreme darkness in the learned global descriptor. Firstly, we establish a day-night urban scene dataset called NightCities, capturing diverse nighttime scenarios and lighting variations across 60 cities globally. Following this, an unpaired image-to-image translation network is trained on this dataset. Using this trained translation network, we process an existing VPR dataset, thereby obtaining its nighttime version. The NocPlace is then fine-tuned using night-style images, the original labels, and descriptors inherited from the Daytime VPR model. Comprehensive experiments on various nighttime VPR test sets reveal that NocPlace considerably surpasses previous state-of-the-art methods.
1 Introduction
Figure 1: A demo for NocPlace. The center of this figure presents a map of San Francisco. 2.8 million daytime database images and 466 nighttime query images from the SF-XL dataset are projected onto this map via GPS tags, corresponding to the dense cyan-blue and sparse red dots, respectively. Despite the dazzling lights and extreme darkness, our method achieves higher recall@N than previous methods.
Visual Place Recognition (VPR) is a foundational task in computer vision [5, 24, 45, 52, 51, 65] and intelligent robotics [20, 32, 34, 36, 39, 58, 33, 61]. The goal of VPR is to retrieve similar database images to a given query image by comparing against a vast set of known database images. Existing literature identifies several challenges for VPR, including database scale [9], viewpoint shifts [8, 11], repeated structures [51], structural modifications [23], occlusions [5], visual scale differences [17], illumination changes [4, 12, 34, 32, 44, 50, 52, 58, 65], and seasonal transitions [36, 49]. Recent VPR approaches primarily leverage large-scale datasets [9, 35, 37, 52, 60, 3], and utilize neural networks to map images into an embedding space, thereby effectively distinguishing images from different locations.
However, night-to-day VPR is constrained by the current training datasets. On the one hand, recent studies [5, 9, 3] leveraging the Google Street View (GSV) download interface have produced large-scale multi-view VPR datasets covering nearly all challenges except those in nighttime conditions. On the other hand, while some datasets specifically targeting night-to-day VPR datasets [34, 36] have been proposed and shown the benefit of incorporating nighttime data for cross-domain recognition, their scale or data acquisition methods impede the training of a universally night-to-day VPR model. It is worth acknowledging that constructing a nighttime urban street-scape dataset aligned with commercial offerings like GSV is challenging.
Given the lack of comprehensive end-to-end night-to-day VPR solutions, image enhancement algorithms (IE) [69, 67, 22, 21] and night-to-day Image-to-Image (I2I) translation [4] may be a potential approach. These methods, however, present several limitations: i) Introducing an auxiliary model significantly increases computational demands. ii) Most low-light enhancement models rely on images with diverse exposure rather than authentic night-to-day transitions, and much of the data is not even sourced outdoors. iii) Imperfect Night-to-Day translations might adversely affect VPR performance [4].
In this paper, we reverse the night-to-day thinking and study how to transfer night information into neural networks directly. Further, we study the use of daytime knowledge to guide the training of nighttime models. Our method can improve the performance of nighttime visual place recognition without consuming additional computing resources.
Our contributions can be summarized as follows:
  • We propose NocPlace, a scalable methodology for Nocturnal Visual Place Recognition, which significantly narrows the VPR performance gap between daytime and nighttime domains.
  • We propose NightCities, an unpaired day-night urban scene image dataset. We utilize this dataset and an I2I translation network to transfer night knowledge to an existing large-scale multi-view daytime VPR dataset.
  • We propose the inherited knowledge transfer which employs descriptors inherited from previous VPR models to guide the training of new VPR models.
  • Extensive experiments demonstrate that NocPlace substantially outperforms previous methods. Specifically, for certain datasets, our VGG16-based model excels over state-of-the-art ResNet50-based models.
  • The source code and trained models for NocPlace are available at TBD.
2 Related Work
Visual Place Recognition can be categorized along three pivotal axes: model architecture, loss function, and dataset. With regard to model architecture, Convolutional Neural Networks (CNNs) [5] and Visual Transformers (ViTs) [53, 56, 10] emerge as the primary choices for feature extraction. These are often followed by pooling or aggregation mechanisms like NetVLAD [5, 41]. In the realm of loss functions, triplet and classification losses are frequently employed to optimize Euclidean margins for enhanced feature embeddings [10, 5]. Ge et al. utilized triplet loss and network output scores as self-supervised indicators, attaining state-of-the-art outcomes via iterative training. However, triplet loss encounters challenges with hard negative sample mining. In response, Berton et al. [9, 11] incorporated Large Margin Cosine (LMC) Loss [55] in VPR, showcasing its superiority over triplet loss. On the dataset front, we observe two primary classifications: self-curated datasets and web-sourced ones. The former, such as [20, 34, 36], often exhibit more minor scales, fixed frontal camera view, and little nighttime data. The latter, however, can be sub-categorized into datasets like Google Landmark [41], derived from Google Images, and those like SF-XL [9] and Pittsburgh [5], sourced from Google Street View [5, 1, 3, 9]. It is worth noting that the latter dataset contains almost no night data. Therefore, it can be inferred that the performance of existing VPR models in nighttime scenarios benefits from data augmentation [10] during the training phase and the generalization capability during the inference phase.
Night Computer Vision encompasses traditional downstream tasks using nighttime imagery, which are primarily divided into two-stage and one-stage approaches. The two-stage approaches [4, 62] first transform nighttime images into their daytime counterparts before executing the downstream tasks. In contrast, one-stage methods directly utilize raw nighttime data for training [14, 65]. For example, Anoosheh et al. [4] introduced a CycleGAN-based method to transform nighttime images for retrieval-based localization. Cui et al. [14] developed a method for reverse ISP and dark processing, leading to an end-to-end model for detecting objects in dark settings. Additionally, Xu et al. [62] presented a GAN-centric strategy for semantic segmentation in autonomous driving scenarios using transformed nighttime imagery. However, within the VPR domain, the two-stage methods are criticized for their limited generalizability [4] and heightened computational demands. Meanwhile, the one-stage approach faces challenges due to the lack of extensive training datasets. In response, we propose a pioneering framework by converting daytime images to nighttime versions for enhanced VPR training.
Image-to-Image Translation pertains to the intersection of vision and graphics, aiming to establish a mapping between an input and an output image through a designated training set. These methods are primarily categorized based on the nature of their training sets: pixel-level matched I2I [27] or unpaired I2I [71, 64]. Securing extensive pixel-level correspondent day-night image pairs in real-world scenarios is notably challenging. Hence, our focus is on the latter training set type. Two predominant strategies for unpaired image-to-image I2I translation exist: bi-directional mapping and uni-directional mapping. Bi-directional mapping, encompassing CycleGAN [71] and DualGAN [64], is anchored on the cycle-consistency constraint, ensuring the transposed image can undergo reverse mapping for reconstruction. The bi-directional mapping can be overly restrictive, prompting the evolution of uni-directional mapping techniques. Some techniques employ geometry distance for content retention. For instance, DistanceGAN [16] conserves inter-image distances within domains, while GCGAN [7] upholds geometry consistency between inputs and outputs. The CUT method [38] amplifies mutual information between input and output via contrastive learning. Recently, the diffusion model [70] has showcased exceptional efficacy in addressing I2I translation challenges, albeit with significant computational demands. After an exhaustive evaluation of various strategies, our day-to-night I2I translation model was developed, rooted in the NEGCUT framework [57].
Knowledge Transfer can be categorized in various ways based on the problems it addresses, such as domain adaptation [66], multi-task learning [68], meta-learning [54], and more. This paper focuses on domain adaptation, aiming to adapt models trained in one domain (source) to perform optimally in another domain (target) where the data distribution might differ. Previous Night-to-Day VPR studies primarily focused on labeled target domains, a relatively straightforward problem [61, 63, 40]. Cao et al. [40] incorporated Canny edge descriptors and the Wasserstein distance measure for domain adaptation in VPR tasks. When dealing with labeled source and unlabeled target domains, prior research can be broadly divided into two categories: i) generation of target domain data, as exemplified by SimGAN [47], which transfers samples from the source to the target domain while preserving consistent labels. ii) adversarial learning of domain-invariant features. Ganin et al. introduced DANN [18], which learns domain-invariant feature distributions by minimizing the source domain classification loss and maximizing the domain classification loss. We present the generative knowledge transfer, which aligns with the first category, and introduce inherited knowledge transfer that minimizes the target domain classification loss and the discrepancy between domain feature distributions.
3 Day-to-Night Translation Datasets
In this section, we elucidate the motivation behind constructing the NightStreets and NightCities datasets. Subsequently, we delve into details regarding the composition of these datasets. It is imperative to note that these two datasets are prerequisites for our methodology.
Figure 2: Potential Datasets for Day-to-Night Translation. (a) LOL dataset [59], (b) RobotCar dataset [34], (c) night2day dataset [29], and (d) NightStreets dataset.
3.1 Motivation for Datasets Construction
We present a novel observation: existing large-scale, multi-view VPR training sets, ranging from 30k [5] to an impressive 41.2M images [9], are predominantly composed of daytime images and lack nighttime representation. This omission significantly affects their recognition capabilities under nocturnal conditions, rendering them sub-optimal compared to daytime scenarios. While some day-night VPR training sets [34, 58] have demonstrated the benefits of nighttime training data for improved cross-domain recognition, they suffer from limited generalization. Two main factors contribute to this limitation: i) a consistent nighttime style across the datasets and ii) a considerably smaller scale compared to daytime datasets. We also recognize two major challenges: i) Constructing a nighttime VPR dataset that aligns with the GPS tags or 6 DoF poses of existing VPR datasets is exceptionally difficult. ii) Building a Day-Night VPR dataset from scratch, which matches the scale of current datasets, is also a daunting task. Given these challenges, we propose an indirect approach: utilizing I2I translation techniques to model daytime images in a nighttime style. Consequently, there is an urgent need for datasets encompassing both day and night scenarios for training an I2I translation network.
Regrettably, our search for a suitable day-to-night I2I translation dataset was unsuccessful. As illustrated in Fig. 2 (a), most low-light enhancement datasets are primarily concerned with low-light conditions [59, 69], such as those induced by backlighting or weak exposure, rather than authentically capturing the diverse nuances of real nighttime scenes. Fig. 2 (b) showcases examples from the aforementioned day-night VPR dataset [34]. Models trained on these datasets have limited generalization capabilities for night-to-day VPR, suggesting that theoretically generated day-to-night scene data might also be inadequate. Furthermore, we observed recurring limitations such as static camera orientations due to vehicular mounting, leading to image blur. Fig. 2 (c) references several time-lapse photography datasets proposed in the previous I2I transfer research [29, 27]. However, these datasets tend to focus on distant views and skylines, deviating significantly from urban scenes, and a day-to-night time-lapse video contains higher redundancy than a few photos.
Figure 3: Examples of the NightCities dataset. Only nighttime images are displayed here. From left to right, top to bottom, they represent Tokyo, Paris, Zurich, Hong Kong, Washington, San Francisco, Sydney, Munich, and Vancouver.
3.2 The NightCities Dataset
Consequently, we collected a large-scale unpaired day-night urban scene image dataset named NightCities, as illustrated in Fig. 3. It encompasses 5,000 images from 60 cities globally, encapsulating diverse lighting styles, cultural vibes, and architectural features. Notably, the quantity of images aligns with representative datasets in the I2I translation domain [13, 28, 70]. The only annotation required was the categorization into day and night. We trained an unpaired I2I translation network on the NightCities dataset and used it to process a subset of the SF-XL dataset, resulting in SF-XL-NC. It consists of a training set of 59,650 night-style images as well as a validation set of 7,993 night-style queries and 8,015 database images.
3.3 The NightStreets Dataset
We also built a test set for evaluating I2I translation by reorganizing query images from Tokyo 24/7 v3 [52] and Aachen Day/Night v1.1 [46], as illustrated in Fig. 2 (d). Tokyo 24/7 v3[52] comprises 375 daytime and nighttime images each, while Aachen Day/Night v1.1 [46] contains 234 daytime and 196 nighttime images. To distinguish it from the previous data sets, we call it NightStreets.
4 Methodology
Figure 4: Schematic Diagrams of NocPlace. In the training phase, we transfer the night knowledge into the training data and use the existing label information. We inherit the descriptors extracted from the corresponding data by the pre-trained daytime VPR model and use them to supervise the NocPlace model. During the test phase, NocPlace extract features from the query image and query them in the descriptor database extracted by the daytime VPR model.
In this section, we detail the training and testing strategies of NocPlace, which encompass generative knowledge transfer, inherited knowledge transfer, and a divide-and-conquer retrieval approach. The upper part of Fig. 4 corresponds to Sec. 4.1 and 4.2, and the lower part corresponds to Sec. 4.3.
4.1 Generative Knowledge Transfer
Generative knowledge transfer aims to generate pseudo target domain data using labeled source domain data, especially in scenarios where labeled data in the target domain is hard to obtain. The labels from the source domain data are then directly transferred to the corresponding generated data. Specifically, our objective is to maintain the scene’s content from daytime images while coupling it with a nighttime style. For implementation, we employed NEG-CUT [57], an unpaired I2I generation method based on contrastive learning [38]. It maintains local content consistency by mining challenging image patches and learns the target domain style using a discriminator and adversarial loss [27].
Employing the VPR training framework [9, 11] with Large Margin Cosine (LMC) loss [55], we can either fine-tune the model using both source and generated domain data or train exclusively with the generated domain data. Theoretically, the former approach balances performance between daytime and nighttime, while the latter emphasizes nighttime performance.
4.2 Inherited Knowledge Transfer
The advantage of LMC loss over triplet loss is that it avoids mining difficult samples, which makes it easier to achieve the current state-of-the-art results on large-scale training data sets [9, 11]. However, LMC loss also has the disadvantage that it cannot explicitly construct cross-domain contrastive relationships. Meanwhile, we assume that a pre-trained VPR model has learned a representation suitable for daytime VPR. Therefore, we hope to minimize the distribution difference between the day and night domains to inherit the representation ability of the pre-trained model. Specifically, we save the global descriptors extracted from daytime images by the pre-trained model, align them with the nighttime images and minimize the network outputs by Kullback–Leibler (KL) divergence.
We call the above method Inherited Knowledge Transfer (IKT). It is very similar to Knowledge Distillation (KD) [26] at a macro level but is significantly different in specific definition and implementation. KD involves using the outputs from a larger model (teacher) to guide the training of a smaller model (student) on the same dataset, aiming to enhance the student’s performance. In most cases, the performance of this student model will not exceed that of the teacher model. In our case, we guide a new model with an older model of the same architecture on data containing new knowledge, aiming to enhance the new model’s performance.
In terms of implementation, IKT loss is also different from KD loss. IKT Loss is aligned with LMC loss and is calculated in cosine space. The formula for LMC loss is as follows:
$L_{\mathrm{LMC}}=\frac{1}{N}\sum_{i}{-\log{\frac{e^{s(\cos(\theta^{(n)}_{{y_{i}},i})-m)}}{e^{s(\cos(\theta^{(n)}_{{y_{i}},i})-m)}+\sum_{j\neq y_{i}}{e^{s\cos(\theta^{(n)}_{j,i})}}}}},$ (1)
subject to
$\begin{split}cos(\theta^{(n)}_{j},i)&={W_{j}}^{T}x^{(n)}_{i},\end{split}$ (2)
where $N$ is the number of training images, $x^{(n)}_{i}$ is the $i$-th embedding vector extracted from night-style data corresponding to the ground truth class of $y_{i}$, the $W_{j}$ is the weight vector of the $j$-th class, and $\theta^{(n)}_{j}$ is the angle between $W_{j}$ and $x^{(n)}_{i}$. $s$ and $m$ are two hyper-parameters that control the weights of the intra-class distance and inter-class distance in the loss function, respectively.
Building on the previous notation, the softened probabilities obtained from the pre-trained model on the i-th daytime data can be expressed as:
$\begin{split}p^{(d)}_{i}=\frac{e^{s(\cos(\theta^{(d)}_{{y_{i}},i})-m)}}{e^{s(\cos(\theta^{(d)}_{{y_{i}},i})-m)}+\sum_{j\neq y_{i}}{e^{s\cos(\theta^{(d)}_{j,i})}}}.\end{split}$ (3)
Incorporating the above expression for softened probabilities $p^{(d)}_{i}$ and $p^{(n)}_{i}$ into the following equation:
$\begin{split}L_{\mathrm{IKT}}=\frac{1}{N}\sum_{i}{D_{\mathrm{KL}}(p^{(d)}_{i}||p^{(n)}_{i})},\end{split}$ (4)
where $D_{\mathrm{KL}}$ represents the KL divergence. Finally, the final loss is:
$\begin{split}L=L_{\mathrm{LMC}}+\alpha L_{\mathrm{IKT}},\end{split}$ (5)
where $\alpha$ is a hyperparameter used to balance the two losses.
4.3 Retrieval based on Partial Divide-and-Conquer
We have decomposed VPR into two distinct sub-problems, Day-to-Day VPR and Night-to-Day VPR, addressing these challenges using two separate models. This approach is novel, and we explore its practicality and potential for optimization.
In real-world applications, computers can easily differentiate between day and night using straightforward methods that require only one-time operations. For instance, we can determine day and night simply by the system time of a device. Therefore, NocPlace involves only model parameter switching within the same architecture, without introducing additional inference overhead. This significantly enhances performance in nighttime environments.
An inherent limitation of NocPlace is the requirement for computers to store two models and two databases. With the advent of an edge-cloud collaborative VPR system and the ease of scalable storage solutions in modern computing, this drawback becomes a manageable trade-off. However, if both the database and the model are on the edge, such as in decentralized autonomous robots, it would significantly consume storage resources. For instance, each image in the SF-XL test set is output as a feature vector of 512 dimensions, and the entire database size would be a staggering 5.7 GB. To overcome this challenge, we only use NocPlace to process the query image during the testing phase, while the database retains the features extracted by the pre-trained model. This design also aligns with the design of IKT.
5 Experiments
In this section, we begin by providing a description of the implementation details, the test sets, and the evaluation metrics. Subsequently, we present a comprehensive range of quantitative experiments focusing on night-to-day visual place recognition, night-to-day visual localization, and day-to-night image-to-image translation.
Dataset Name SVOX SF-XL MSLS Tokyo 24/7 Tokyo 24/7 Aachen v1.1 RobotCar Seasons v2 Night Night Night Night Sunset Night Night # queries 823 466 55 105 105 192 438 # database 17k 2.8M 18.9k 76k 76k 6.7k 6.9k Orientation frontal-view multi-view frontal-view multi-view multi-view multi-view frontal-view Scenery mostly urban urban urban urban urban mostly urban mostly urban Acquisition Method vehicles handheld devices vehicles handheld devices handheld devices handheld devices vehicles
Table 1: Overview of night VPR and VL test sets. We can see that the test sets listed above includes various scales, scenarios, collection methods, and database image orientations.
5.1 Implementation details
5.1.1 Night-to-Day Visual Place Recognition
Architecture. The core concept of NocPlace revolves around transferring generated nocturnal information and existing day information into a new model. Therefore, it is expected to be compatible with all previous methods and models, enhancing their performance specifically in nocturnal scenarios. In order to illustrate the potential and scalability of NocPlace, we followed a methodology akin to CosPlace [9] and EigenPlaces [11], leveraging a simple convolutional neural network architecture, specifically VGG-16 [48] and ResNet-50 [25]. The architecture was connected to a GeM pooling layer [41] and a fully connected layer, with the dimensionality of the descriptors produced by NocPlace being determined by the specifications of the fully connected layer.
Training. NocPlace employed CosPlace (512 dimensions) or EigenPlaces (2048 dimensions) as pre-trained models, requiring only 1 epoch of fine-tuning on SF-XL-NC with a learning rate of 1e-6. An important distinction is that NocPlace no longer utilizes any data augmentation techniques. The parameter $\alpha$ in the loss function can be adjusted by observing the initial loss of $L_{\mathrm{LMC}}$, but it does not require meticulous tuning. For instance, in training a model based on the ResNet-50 with an output dimension of 512, this parameter was set to 30.
5.1.2 Night-to-Day Visual Localization
We adopt the hierarchical localization framework proposed by [42], utilizing Superpoint [43] for local feature extraction and LightGLUE [15, 30] for feature matching. The key modification we introduce lies in the VPR module, where we adjust the input image resolution to align it appropriately with the framework’s predefined parameters.
5.1.3 Day-to-Night Image-to-Image Translation
We did not modify any training parameters and followed the official tutorial directly [38, 57]. It is worth noting the difference in image resolution between NightCities and the VPR training sets. To address this, we proportionally resize the NightCities images, keeping the shorter side at 640 pixels, and subsequently apply random cropping to achieve a size of 512$\times$512 during the training process. This approach ensures as much resolution consistency as possible between the I2I and VPR training sets. During the evaluation, we trained NEGCUT on multiple data sets mentioned in Sec. 5 and used NightStreets as the test set.
5.2 Test sets
We conducted experiments on multiple (6) nighttime VPR datasets, which encompass challenges frequently discussed in daytime VPR scenarios: large-scale, viewpoint variations, scale changes, etc. Some specific details are presented in Tab. 1. Aside from adaptive modifications to the Tokyo 24/7 [5] and MSLS [58] datasets, the rest follow prior works [45, 65, 46, 11].
Tokyo 24/7 [5] is highly challenging in terms of scale, with a vast database size and a three-year gap between query images and database images. Notably, the query images for each location are distributed across three time periods: daytime, sunset, and nighttime. To the best of our knowledge, few studies [6, 5] segment this test set according to these time frames, somewhat overshadowing the impact of nighttime.
MSLS [58] is a crowdsourced dataset derived from dashcam recordings. We followed the validation set configuration from [9] and filtered out nighttime images. Although our experiment approach differs from [11] and other studies, it ensures a fair comparison in our experiments.
Method Backbone
Feat.
Dim.
Training
Dataset
SVOX
Night
SF-XL
Night
MSLS
Night
Tokyo 24/7
Night
Tokyo 24/7
Sunset
NetVLAD [5] (CVPR’16)
SFRS [19] (ECCV’20)
CAE-VPR [63] (Robotica’23)
CosPlace [9] (CVPR’22)
EigenPlaces [11](ICCV’23)
NocPlace (Ours)
VGG-16
VGG-16
VGG-16
VGG-16
VGG-16
VGG-16
4096
4096
4096
512
512
512
Pitts-30k
Pitts-30k
RobotCar
SF-XL
SF-XL
SF-XL-NC
8.0 / 17.4 / 23.1
28.7 / 40.6 / 46.4
1.1 / 4.1 / 5.8
44.8 / 63.5 / 70.0
42.3 / 61.0 / 68.5
53.0 / 71.2 / 76.4
6.7 / 12.0 / 14.2
11.6 / 18.9 / 22.1
1.3 / 2.4 / 3.0
17.8 / 24.7 / 28.3
18.5 / 26.4 / 28.1
20.8 / 28.8 / 31.8
1.8 / 1.8 / 3.6
1.8 / 7.3 / 14.5
1.8 / 1.8 / 1.8
10.9 / 16.4 / 29.1
1.8 / 12.7 / 18.2
29.1 / 50.9 / 56.4
38.1 / 58.1 / 61.0
61.0 / 77.1 / 84.8
14.3 / 28.6 / 34.3
65.7 / 84.8 / 87.6
67.6 / 81.0 / 84.8
79.0 / 86.7 / 91.4
81.9 / 89.5 / 90.5
87.6 / 91.4 / 91.4
21.9 / 32.4 / 38.1
89.5 / 90.5 / 93.3
87.6 / 94.3 / 97.1
88.6 / 91.4 / 94.3
CosPlace [9](CVPR’22)
Conv-AP [1] (NC’22)
MixVPR [2] (WACV’23)
EigenPlaces [11] (ICCV’23)
NocPlace (Ours)
ResNet-50
ResNet-50
ResNet-50
ResNet-50
ResNet-50
512
512
512
512
512
SF-XL
GSV
GSV
SF-XL
SF-XL-NC
51.6 / 68.8 / 76.1
36.0 / 52.5 / 61.2
44.8 / 63.2 / 71.0
51.5 / 70.8 / 78.4
63.9 / 77.9 / 84.9
23.8 / 29.0 / 31.5
4.5 / 7.9 / 10.1
10.7 / 18.2 / 20.0
22.3 / 29.8 / 32.8
26.6 / 32.0 / 34.5
7.3 / 10.9 / 16.4
1.8 / 9.1 / 10.9
1.8 / 12.7 / 20.0
3.6 / 10.9 / 12.7
65.5 / 74.5 / 85.5
80.0 / 88.6 / 91.4
39.0 / 58.1 / 65.7
55.2 / 71.4 / 76.2
79.0 / 87.6 / 90.5
89.5 / 93.3 / 94.3
91.4 / 97.1 / 99.0
67.6 / 80.0 / 88.6
82.9 / 93.3 / 95.2
95.2 / 98.1 / 98.1
94.3 / 96.2 / 98.1
CosPlace [9](CVPR’22)
Conv-AP [1] (NC’22)
EigenPlaces [11] (ICCV’23)
[11] + ToDay-GAN [4] (ICRA’19)
NocPlace (Ours)
ResNet-50
ResNet-50
ResNet-50
ResNet-50
ResNet-50
2048
2048
2048
2048
2048
SF-XL
GSV
SF-XL
RobotCar
SF-XL-NC
50.8 / 67.6 / 74.8
37.9 / 57.1 / 65.4
58.8 / 76.9 / 82.6
13.0 / 27.9 / 36.0
75.6 / 87.5 / 91.4
23.6 / 29.2 / 32.8
5.6 / 10.1 / 12.0
23.6 / 30.9 / 34.5
3.9 / 8.6 / 10.9
28.3 / 35.8 / 39.7
5.5 / 10.9 / 16.4
1.8 / 9.1 / 18.2
3.6 / 10.9 / 10.9
12.7 / 20.0 / 25.5
69.1 / 76.4 / 80.0
78.1 / 89.5 / 91.4
47.6 / 60.0 / 66.7
83.8 / 92.4 / 94.3
22.9 / 40.1 / 48.7
91.4 / 94.3 / 96.2
89.5 / 94.3 / 96.2
77.1 / 86.7 / 89.5
96.2 / 98.1 / 98.1
8.6 / 17.2 / 24.8
93.3 / 99.0 / 99.0
Table 2: Recall@1 / 5 / 10 on Night-to-Day VPR datasets. The best results under the same model and the same data set are bolded.
Method Backbone Feat. Dim. Aachen v1.1 Night RobotCar Seasons v2 Night
Recall@1 Recall@5 Recall@10 Recall@1 Recall@5 Recall@10
NetVLAD[5]
SFRS[19]
CAE-VPR [63]
Conv-AP [1]
CosPlace [9]
EigenPlaces [11]
[11] + ToDay-GAN [4]
NocPlace (Ours)
VGG16
VGG16
VGG16
ResNet-50
ResNet-50
ResNet-50
ResNet-50
ResNet-50
4096
4096
4096
2048
2048
2048
2048
2048
63.4 / 80.6 / 88.0
69.6 / 85.3 / 95.3
46.6 / 55.5 / 64.4
57.6 / 77.0 / 85.3
66.5 / 84.3 / 95.8
67.5 / 83.2 / 93.7
22.5 / 31.4 / 38.2
68.6 / 85.9 / 95.8
68.6 / 87.4 / 97.4
71.2 / 87.4 / 97.4
57.6 / 70.2 / 78.5
71.2 / 85.9 / 94.8
71.2 / 86.9 / 97.9
72.8 / 88.5 / 97.9
37.7 / 50.3 / 57.6
72.8 / 89.0 / 96.9
70.7 / 85.9 / 95.3
71.7 / 88.0 / 98.4
60.7 / 77.0 / 83.8
74.9 / 89.0 / 96.9
73.8 / 88.0 / 97.9
73.8 / 88.0 / 97.9
42.9 / 56.0 / 64.9
74.9 / 90.1 / 99.0
11.5 / 22.1 / 30.1
18.1 / 35.8 / 48.2
25.7 / 55.8 / 67.7
21.2 / 38.5 / 51.3
14.6 / 26.1 / 40.7
14.2 / 23.9 / 38.5
14.6 / 25.2 / 40.7
22.1 / 41.6 / 58.8
18.1 / 38.9 / 50.9
21.7 / 50.0 / 63.7
34.1 / 69.5 / 81.0
25.7 / 56.2 / 67.3
19.0 / 34.1 / 48.2
16.8 / 31.0 / 46.0
19.9 / 39.4 / 54.0
23.9 / 54.4 / 72.6
21.2 / 46.9 / 58.0
23.5 / 59.3 / 70.8
32.7 / 72.6 / 84.1
29.2 / 62.4 / 75.7
19.0 / 36.3 / 51.3
18.1 / 36.7 / 50.4
27.4 / 48.2 / 61.5
28.8 / 58.8 / 79.2
Table 3: Localization success rate at different localization thresholds, recall@N, and datasets. [50]. The localization thresholds are (0.25m, 2°) / (0.5m, 5°) / (5m, 10°). Specifically, recall numbers are set to 1, 5, 10. The best results for the same data set and the same threshold are bolded.
5.3 Evaluation metrics
Night-to-Day Visual Place Recognition. For VPR datasets, our evaluation metric is recall@N using a 25-meter criterion. Specifically, this denotes the fraction of query instances where at least one among the top N predictions falls within a 25-meter proximity to the query’s ground truth location, as commonly adopted in prior studies [5, 52, 39, 9, 31, 19].
Night-to-Day Visual Localization. We evaluate VPR through the localization success rate under varying recall@N metrics. It is crucial to note that the visual localization task demands more rigorous precision than conventional VPR benchmarks.
Day-to-Night Image-to-Image Translation. Our evaluation of translated images encompasses two primary facets: realism and faithfulness [71, 38, 57]. Realism quantitatively assesses the disparity between the target and translated image distributions. Faithfulness gauges the extent of congruence between an original image and its translation on an individual basis. For a precise measure of realism, we leverage the established Frechet Inception Score (FID). In measuring Faithfulness, we utilize metrics like the pixel-wise $L_{2}$ distance, peak-signal-to-noise ratio (PSNR), and the structural similarity index measure (SSIM) for each paired input-output.
5.4 Comparison with previous work
We have conducted extensive comparisons of numerous methods on various datasets, an endeavor unprecedented in nighttime VPR, as shown in Tab. 2 and Tab. 3. These methods range from the classical NetVLAD[5], To-Day GAN [4] to more recent contributions over the past three years, including SFRS [19], CosPlace [9], MixVPR [2], CAE-VPR [63], Conv-AP [3], and EigenPlaces [11]. We replicated their open-source models and code for our experiments. It is worth noting that these methods were also tested on the SVOX Night dataset in [11] and our results are consistent with theirs. The experimental results can be distilled into the following key observations:
  • NocPlace is the undisputed winner in nighttime scenarios. Its performance enhancements are substantial, even when compared to EigenPlaces (ICCV’2023).
  • The sunset scenes in Tokyo 24/7 present a mix of numerous daytime and fewer nighttime images, which exposes the degradation of NocPlace during daylight conditions.
  • On both SVOX Night and Tokyo 24/7 Night datasets, NocPlace achieves a significant improvement in cross-architectural performance.
  • The advantages of NocPlace are magnified on the MSLS Night dataset, due to the limited number of query images (less than 100) and redundancy (sequential frames).
  • In scenarios with particularly large databases, such as the SF-XL dataset, the improvements of NocPlace are objectively manifested. Conversely, in smaller databases like the Aachen v1.1 dataset where the VPR challenge is reduced, the gains of NocPlace become less pronounced, even trailing behind methods with higher feature dimensions, such as SFRS.
  • On the RobotCar Seasons v2 Night dataset, CAE-VPR achieved the best results, and ToDay GAN was able to enhance the performance of EigenPlaces. However, they showed poor generalization on other datasets, which confirms the viewpoint we stated earlier. Additionally, it is noteworthy that the training set chosen by CAE-VPR and the test set for VL seem to have an intersection (Their division methods are inconsistent).
5.5 Ablations
Training set IKT OD Tokyo 24/7 SVOX Day Night Night Night 80.0 / 88.6 / 92.4 51.6 / 68.8 / 76.1 83.8 / 94.3 / 94.3 62.3 / 77.4 / 81.9 83.8 / 93.3 / 96.2 63.2 / 78.7 / 84.8 89.5 / 93.3 / 94.3 63.9 / 77.9 / 84.9 86.7 / 92.4 / 92.4 65.6 / 80.8 / 85.3 87.6 / 93.3 / 92.4 66.6 / 80.9 / 87.1 91.4 / 93.3 / 93.3 67.3 / 81.0 / 86.6
Table 4: Ablation on designs, shown as the Recall@ 1/5/10 obtained with a ResNet-50 with output dimensionality 512 on two datasets.
We conducted ablation experiments on the ResNet-50 architecture with an output dimension of 512 for NocPlace, as shown in Tab. 4. It is evident that generating nighttime data and incorporating the IKT loss both result in significant performance improvements. Here, OD means using the original database, which also means the partial divide-and-conquer based retrieval mentioned before. By using OD, we sacrificed accuracy to maintain the practicality of the method. This implies that without using OD, NocPlace in Tab. 2 and Tab. 3 could achieve even better experimental results.
5.6 Comparision with previous datasets
Dataset FID $\downarrow$ L2 $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ Night2Day 243.90 164.79 8.74 0.21 LOL 99.62 191.99 7.40 0.14 RobotCar 148.88 156.29 9.29 0.32 NightStreets 63.54 121.62 11.75 0.60 NightCities 62.21 98.12 13.98 0.75
Table 5: Quantitative comparison of generated results for different Night-to-Day datasets. The best results are shown in bold.
We focus on the quality of the day-to-night dataset. To provide an objective assessment regarding the dataset in our work, we trained NEG-CUT on the datasets previously mentioned and on the dataset we proposed. Only the Night2Day dataset uses the pre-trained model of Pix2Pix [27] because it is a paired day-night dataset. The test set directly utilized NightStreets. As shown in the Tab. 5, the NightCities model achieved the best results on evaluation metrics commonly used in the I2I translation domain, even surpassing the NightStreets model.
6 Conclusion
In this paper, we introduce a novel training and testing strategy tailored for VPR, aimed at tackling the challenges of cross-domain recognition between day and night scenarios. With the proposed unpaired day-night location dataset, we can transfer the appearance knowledge of nighttime to existing large-scale, multi-view VPR datasets. By minimizing the classification loss of the target domain and the domain discrepancy between the source and target domains, we further enhance the performance of night-to-day VPR. Extensive experiments have been conducted on comprehensive nighttime VPR datasets to substantiate our contributions. NocPlace, with its straightforward and scalable approach, not only addresses the degradation observed in existing VPRs during night-to-day recognition but will also benefit future VPR systems.
Supplementary Material
7 Datasets
7.1 NightCities
We supplement the NightCities with daytime instances and category directories, as shown in Fig. 5 and Fig. 6. Over the next 1-3 years, we will continue to maintain and expand this dataset.
Figure 5: Daytime examples of the NightCities dataset. We supplement the daytime examples of the NightCities dataset here. From left to right, top to bottom, they represent Tokyo, Paris, Zurich, Hong Kong, Washington, San Francisco, Sydney, Munich, and Vancouver.
Figure 6: Histogram of the number of photos for each city. We have collected daytime and nighttime street view images of 60 cities and verified their accuracy manually. Over the next 1-3 years, we will continue to maintain and expand this dataset.
7.2 SF-XL-NC
We trained NEG-CUT on the NightCities dataset and used it to process a subset of SF-XL, resulting in SF-XL-NC. As shown in Fig. 7, we present some examples demonstrating our success in coupling nighttime knowledge with other desired knowledge for VPR.
Figure 7: Partial instances of the SF-XL-NC dataset. The second, fourth, and sixth rows are images from three locations in SF-XL, while the first, third, and fifth rows are the corresponding night dataset generated for the aforementioned three rows.
8 Experiments
8.1 Comparing I2I translation methods
Method FID $\downarrow$ L2 $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ Avg Time $\downarrow$ CycleGAN 63.93 117.59 12.88 0.68 10 ms CUT 63.20 129.47 11.12 0.60 3.2 ms NEGCUT 63.54 121.62 11.75 0.60 3.2 ms ESGDE - - - - 62000 ms
Table 6: Comparsion of different unpaired I2I methods
We trained a range of state-of-the-art unpaired I2I generation methods to the NightStreets dataset, as detailed in Tab. 6. Interestingly, the disparities in outcomes among these methods are considerably less pronounced than those observed on other public datasets, presenting a deviation from some established research trends. However, one aspect is clear: the increasingly popular Diffusion Model, though capable of generating high-quality images, is notably time-intensive. For example, processing a training set of 100,000 VPR images with ESGDE [70] on a 3090 Ti GPU requires approximately 72 days. Given that our primary research focus remains on VPR, we did not extensively dwell on these issues and adopted NEG-CUT [57] for our study.
8.2 Qualitative results
Some qualitative results are shown in Fig 8 and Fig. 9. The figure allows to understand the strengths of NocPlace in a more visual and intuitive way.
Figure 8: Qualitative results on the SF-XL Night dataset. Each column represents a query (first and fourth rows) and the first predicted result from the database. We can see that NocPlace handles extreme lighting variations better than previous methods.
Figure 9: Qualitative results on the Tokyo 24/7 dataset. Each column represents a query (first rows) and the first predicted result from the database. We can see that NocPlace handles extreme lighting variations better than previous methods.